How would you slove this word problem?
A plane traveled 1080 miles to San Francisco and back. The trip there was with the wind. It took 10 hours. The trip back was into wind. The trip back took 15 hours. What is the speed of the plane in still air? What is the speed of the wind?

Answer by scott8148(6628)

rate = distance / time

p + w = 1080 / 10 = 108

p - w = 1080 / 15 = 72

adding equations ___ 2p = 180 ___ p = 90

substituting ___ 90 + w = 108 ___ w = 18